# RadMamba: Efficient Human Activity Recognition through Radar-based Micro-Doppler-Oriented Mamba State-Space Model

### v1

@@ VII-C1 Projection Replacement
The adoption of a 1D convolutional layer as the projection mechanism consistently outperforms configurations with a single linear layer or three linear layers across nearly all settings. Comparing Rows 19-27 with Rows 10-18 and Rows 1-9, the conv1d projection replacement within the same patch size and down-sampling scale achieved average increment of mean accuracy by 0.48% over 1 linear layer, 0.40% over 3 linear layer on Dataset DIAT. On Dataset CI4R and UoG2020, the conv1d layer shows a substantial improvement of 2.99% and 3.64% over 1 linear layer, respectively.
VII-C2 Patch Size
@@
The ($H_{c-d}$, 1) configuration consistently delivers superior performance within each projection layer type. For the 1D convolutional layer, Row 25 with ($H_{c-d}$, 1) and without down-sampling achieves 3% gain on Dataset DIAT, 22.4% gain on Dataset CI4R, 9.2% gain on Dataset UoG2020 vs. Row 19 with rectangular patch. This trend holds across other projection types, indicating that the ($H_{c-d}$, 1) patch size effectively retains the critical spatial features for radar data processing.
VII-C3 Down-sampling
@@
The results suggest that down-sampling along the patch vector dimension generally outperforms down-sampling along the dimension orthogonal to patch vector and no down-sampling. For instance, Row 27’s (2, 32) time down-sampling achieves 88.6% $\pm$ 1.46, outperforming Row 26’s (8, 2) at 84.4% $\pm$ 2.05 and Row 25’s (1, 1) at 84.5% $\pm$ 0.77 on Dataset UoG2020. This pattern highlights the advantage of downsampling in enhancing the micro-Doppler features.
@@
The ablation study underscores the efficacy of the proposed RadMamba design, with the highest performance achieved by integrating a 1D convolutional projection layer, a ($H_{c-d}$, 1) patch size, and time-dimension down-sampling (Row 27).
VIII Conclusion
@@
RadMamba sets a new benchmark in RadHAR, delivering top-tier accuracy with minimal parameters and computational cost across diverse datasets. Its radar micro-Doppler-oriented design addresses domain-specific challenges, making it ideal for real-time, resource-constrained applications. Overall, RadMamba consistently outperforms previous neural network models like ResNet, Bi-LSTM, CNN-LSTM, and CNN-Bi-GRU, and competes with or exceeds advanced architectures like Vision Transformers and ActivityMamba, all while maintaining a remarkably low parameter count (ranging from 6.7 k to 71.4 k) and reduced computational complexity (1.3 M to 145.6 M ##\##FLOP/Inf.). These results emphasize RadMamba’s capability as a lightweight and high-performance solution for radar-based HAR across diverse radar modalities and activity recognition tasks. The micro-Doppler-oriented design, integrating channel fusion, Doppler-aligned segmentation, and convolutional projections, addresses the unique challenges of radar data, delivering high accuracy across CW and FMCW datasets, both continuous and non-continuous. Ablation studies confirm the synergy of its components, with no single element achieving comparable performance alone.
@@ Acknowledgments
This work was partially supported by the European Research Executive Agency (REA) under the Marie Skłodowska-Curie Actions (MSCA) Postdoctoral Fellowship program, Grant No. 101107534 (AIRHAR). We thank Prof. Dr. L.C.N. de Vreede and Prof. Dr. Alexander Yarovoy for their support.

### v3

@@ VI-C1 Projection Replacement
The adoption of a 1D convolutional layer as the projection mechanism consistently outperforms configurations with a single linear layer or three linear layers across nearly all settings. Comparing Rows 19-27 with Rows 10-18 and Rows 1-9, the conv1d projection replacement within the same patch size and downsampling scale achieved an average increment of mean accuracy by 0.48% over 1 linear layer, 0.40% over 3 linear layers on Dataset DIAT. On datasets CI4R and UoG20, the conv1d layer shows a substantial improvement of 2.99% and 3.64% over 1 linear layer, respectively.
VI-C2 Patch Size
@@
The ($H_{c-d}$, 1) configuration consistently delivers superior performance within each projection layer type. For the 1D convolutional layer, Row 25 with ($H_{c-d}$, 1) and without downsampling achieves 3% gain on Dataset DIAT, 22.4% gain on Dataset CI4R, 9.2% gain on Dataset UoG20 vs. Row 19 with rectangular patch. This trend holds across other projection types, indicating that the ($H_{c-d}$, 1) patch size effectively retains the critical spatial features for radar data processing.
VI-C3 Downsampling
@@
The results suggest that downsampling along the patch vector dimension generally outperforms downsampling along the dimension orthogonal to the patch vector and no downsampling. For instance, Row 27’s (2, 32) time downsampling achieves 88.6% $\pm$ 1.46, outperforming Row 26’s (8, 2) at 84.4% $\pm$ 2.05 and Row 25’s (1, 1) at 84.5% $\pm$ 0.77 on Dataset UoG20. This pattern highlights the advantage of downsampling in enhancing the micro-Doppler features.
@@
The ablation study underscores the efficacy of the proposed RadMamba design, with the highest performance achieved by integrating a 1D convolutional projection layer, a ($H_{c-d}$, 1) patch size, and time-dimension downsampling (Row 27).
VII Conclusion
@@
RadMamba sets a new benchmark in RadHAR (Code: https://github.com/lab-emi/AIRHAR), delivering top-tier accuracy with minimal parameters and computational cost across diverse datasets. Its radar micro-Doppler-oriented design addresses domain-specific challenges, making it ideal for real-time, resource-constrained applications. Overall, RadMamba consistently outperforms previous neural network models like ResNet6, Bi-LSTM, CNN-LSTM, and CNN-Bi-GRU, and competes with or exceeds advanced architectures like Vision Transformers and ActivityMamba, all while maintaining a remarkably low parameter count (ranging from 6.7 k to 71.4 k) and reduced computational complexity (1.3 M to 145.6 M #\#FLOP/Inf.). These results emphasize RadMamba’s capability as a lightweight and high-performance solution for radar-based HAR across diverse radar modalities and activity recognition tasks. The micro-Doppler-oriented design, which integrates channel fusion, Doppler-aligned segmentation, and convolutional projections, addresses the unique challenges of radar data, delivering to deliver high accuracy across both CW and FMCW datasets, both including continuous and non-continuous data. Ablation studies confirm the synergy of its components, with no single element achieving comparable performance on its own.
@@ Acknowledgments
This work is funded by the European Research Executive Agency (REA) under the Marie Skłodowska-Curie Actions (MSCA) Postdoctoral Fellowship program, Grant No. 101107534 (AIRHAR). We thank Prof. Dr. L.C.N. de Vreede and Prof. Dr. Alexander Yarovoy for their support.